Consider the probability distribution of the first year cash flows for the DEF Project: Project Cash...

  1. Consider the probability distribution of the first year cash flows for the DEF Project:

Project Cash Flow

Probability

$1,000

20%

$2,000

60%

$3,000

20%

Please show all work, step by step, for a thumbs up! Thanks!!

  1. Calculate the range of possible cash flows
  2. Calculate the expected cash flow
  3. Calculate the standard deviation of the possible cash flows
  4. Calculate the coefficient of variation of the possible cash flows
Answer #1

a. Range of possible cash flow is calculated using difference between minimum and maximum value of project cash flow .

Here minimum cash flow = $1000 and maximum cash flow = $ 3000 . Thus Range = $ 3000 -$1000 = $2000

b. Expected cash flow = = $1000* 20% + $2000 * 60% + $3000 * 20% =

$ 2000

c. Standard Deviation =

=

= 632.455

Where = expected project cash flow x= project cash flow P(x) = probability of respective project cash flow

d. Coefficient of variation = standard deviation / mean () = 632.455 / 2000 = 0.3162